Status Report

ISS On-Orbit Status 25 Jul 2002

By SpaceRef Editor
July 25, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previou
sly or below.

CDR Valery Korzun was scheduled today for a training session with the PUMA
optical collimator sighting device to perform attitude data correction of the SM’s BINS strap-down guidance system.  [For the BINS correction, the current XPOP Momentum Management flight mode was to be commanded by MCC-H
to Attitude Hold on CMGs from 9:50-10:10am EDT, and returned to Momentum Management when complete.]

FE-1 Peggy Whitson activated the MSG (microgravity science glovebox) and continued SUBSA (solidification using a baffle in sealed ampoules) trouble
shooting.  [Because the SUBSA thermal chamber
did not attain its maximum temperature of 845B0C on 7/23, the SUBSA activities for today were replaced with a SUBSA Eurotherm controller troubleshooting
test of about 35 min. duration.  The purpose was to verify that the thermal chamber can heat-up to its maximum temperature after the ground remotely reconfigures the PCM Eurotherm controller set-point limit.  Instructions were uplinked which directed Peggy through the power up cycle in the correct sequence. The maximum heater temperature for the run was the normal 845
B0C. Peggy did not install a sample for this run, so the sample temperatures
on the MLC read 1.0.]

FE-2 Sergei Treschev unstowed the Russian Plasma Crystal 3 experiment module and connected it to the vacuum line, then ran a pressurization
test and vacuum check.

Korzun worked another session with the EVARM (EVA radiation measurement) payload, first activating the EVARM reader, followed by taking background radiation measurements, later transferring the data to storage and stowing the equipment.

Treschev completed an SLM sound meter calibration test, then took another set of noise level readings in the SM.  Later in the day he transferred
the data to the MEC (medical equipment computer) for downlink to the ground.

Peggy Whitson set up the MEC with the EARQ audiometer equipment. Then, she,
Korzun and Treschev recorded audiograms for about 20 min during their respective activities..

FE-1 Whitson was thanked for her recent soybean plant
photos of the ADVASC (advanced astroculture) experiment.  She received
an additional request for ADVASC photos per today92s task list.  [From the 7/23 photos, the ADVASC team could not conclusively identify the developmental stage of the flowers and pod filling.  When she performs the plant photo activity, she is to use the manual focus mode on the
camera to take pictures of individual pods and flowers.]

CDR Korzun worked on recovering (repriming) the BZh fluids unit #3 of the E
lektron oxygen generator, in consultation with a ground specialist.

Korzun also conducted a crosscheck of several pressure gauges (MV) used in
the SM, DC-1, Soyuz and Lab.

The crew reviewed the procedures for the upcoming EPO (educational payload
operations) activities.

They also completed a one-hour emergency operations drill in case of fire.

In tracking down the anomalous noise in the TVIS treadmill, the crew provided some answers, from which engineers are able to focus more on a mechanical
problem in the TVIS chassis (instead of the transfer case, flywheel or motor box) as the most likely source of the noise. However, there are additional
questions and actions needed before specific components of the TVIS chassis
can be identified as the definite noise source.

The CDRA (carbon dioxide removal assembly) failed to start normally following the bed #1 removal and replacement (R&R) activity last week.  Indications are that there is leakage somewhere in the side of the system containing bed #1. [The signature is similar to what was seen prior to the component R&R.  In order to verify the integrity of bed #2, the CDRA
was activated in single bed mode and has been operating normally in this mode.  MCC-H deactivated the CDRA today, in order to gather and analyze the data and then create a troubleshooting plan in the next week or two

At 11:05 am EDT today, the crew was scheduled to participate in two aud
io/video media events.  One interview will be with KCRG-TV, Cedar Rapid
s, IA, the other with Creston, IA Daily News.

MCC-H/PAO was delighted with yesterday’s downlink message for the Marshall
Space Flight Center and the science interview for the Orange County Register
.  The pure discussion of science was very well received and will be re
played frequently on NASA TV to demonstrate the true purpose of a permanent
human presence in space.  "Many thanks from PAO".

All crewmembers performed their daily physical exercise program.

SpaceRef staff editor.